Grade 11Physical Chemistry

0.54 gm of metal M yields 1.02 gm of its oxide M2O3. The at.omic weight of metal M is??

Dhruv Agja
7 Years agoGrade 11

1 Answer

sahil agrawal
7 Years ago
Let the mass of M be x. so mass of M2O3=2x+48
now write reaction
4M+3O2=2M2O3. 
so 2 moles of M will make 1 mole of M2O3
now moles of M given are 0.56/x and that of M2O3 are 1.02/(2x+48)
So we know that moles of M should be double of moles of M2O3 
therefore (0.56/x)=2(1.02/(2x+48))
giving us X =24..